-1 2/7m=6
what is the answer

Answer by fractalier(6550)

Convert the mixed fraction to improper first...we get
-1 2/7m = 6
(-9/7)x = 6
now multiply both sides by -7/9
x = 6(-7/9)
x = -42/9 now reduce
x = -14/3